OpenStudy (anonymous):

help please. medal for answer. ! find the solution to the system of equations by using either graphing or substitution. y = 2x - 1 and y = x + 3 a. 4,7 b. 7,4 c. -7, -4 d. infinite solutions

OpenStudy (anonymous):

So first, we can immediately eliminate D. In order for it to be an infinite solution, the solution must be true for every value of x. So if we decide x = 1, we would have: 2(1) - 1 and (1) + 3 2 - 1 and 1 + 3 1 and 4 Does that part make sense?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Would you like to use substitution here?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

So, the y's are by itself now. We know that y is equal to 2x - 1, and we know that y is also equal to x + 3 Because these are both equal to y, they are equal to each other. 2x - 1 = x + 3

OpenStudy (anonymous):

So now you have to use Algebra to make x by itself. What is x here?

OpenStudy (anonymous):

sorry, im lost. i got the first part. just lost now.

OpenStudy (anonymous):

Can we say that 7-4 and 6-3 will both equal 3?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

So would you also agree that 7-4 = 6-3? Because they are both equal to 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

That same thing is going to apply to your problem. The y's will both have the same value, because only one value can be applied to y. 2x-1 = x+3 because they will both yield the same value, just like my example above. Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So now, we need x on one side, and a numerical value on the other.  What will x equal here?